OpenStudy (anonymous):

.10(y-1)+.14y=.04y-.3

OpenStudy (bahrom7893):

10(y-1)+14y=4y-30

OpenStudy (bahrom7893):

does it look better now?

OpenStudy (bahrom7893):

10y-10+14y=4y-30 24y-4y=10-30 20y=-20 y=-1

OpenStudy (anonymous):

how about 6(x-2)/2=3(x+6)/9